OpenStudy (anonymous):

a+s=150 45a+28.50s=4770 How do I solve this? help please!!!

OpenStudy (hoblos):

a+s=150 ----> a=150-s replace this in the second equation and solve for s

OpenStudy (anonymous):

follow @hoblos and then in ur long equation do: 45(150-s) +28.50s=4770.

OpenStudy (cwrw238):

you can use elimination method by first multiplying ist equation by 45: 45a + 45s = 6750 45a + 28.5s = 4770 subtract to eliminate a

OpenStudy (anonymous):

consider a= 150 - s substitute the value of a in the second eqn 45 (150- s) + 28.5s =4770 6750 - 45s + 28.5s=4770 -16.5s =-1980 s =120 substitute the value of c in any eqn take a + s =150 a + 120=150 a =30 therefore a =30 and s=120

OpenStudy (anonymous):

thanks a bunch